### 07 Mar 2026 v2

# On solutions of singular Sylvester equations in quaternions

Abstract.
The quaternionic equations axxb=0{a}{x}-{x}{b}=0 and axxb=c{a}{x}-{x}{b}=c are investigated, which are called homogeneous and inhomogeneous Sylvester equations, respectively. Conditions for the existence of solutions are provided. In addition, the general and nonzero solutions to these equations are derived applying quaternion square roots.
1. Introduction
A general linear function of a quaternion variable x{x} has the form [1]
$f_{n}({x})={a}_{1}{x}{b}_{1}+{a}_{2}{x}{b}_{2}+\dots+{a}_{n}{x}{b}_{n},$
where the coefficients ${a}_{j},{b}_{j}$ for $j=1,2,\dots,n$ are nonzero quaternions. Due to the noncommutativity of the quaternion product, $f_{n}$ is in general not homogeneous and hence not linear. However, $f_{n}$ is linear over the real numbers [1]. For $n=2$ and without loss of generality $f_{n}$ may be rewritten as
$f({x})={a}{x}-{x}{b}.$
The function $f$ is called singular if there exists a nonzero quaternion x{x} for which f(x)=0f({x})=0, otherwise it is called regular. Given a quaternion c{c} the equation
$f({x})={a}{x}-{x}{b}={c},$
is referred to as the Sylvester equation. The equation is called homogeneous if c=0{c}=0 and inhomogeneous otherwise. The Sylvester equation is called singular if the associated function $f$ is singular, otherwise regular.
Johnson [2] distinguished between the regular and singular cases and provided topological proofs for the existence and uniqueness of solutions over division rings, which directly apply to quaternions. A regular Sylvester equation has a unique solution for any quaternion c{c}. Hence, the homogeneous regular Sylvester equation has only the trivial solution x=0{x}=0. Closed-form expressions for the inhomogeneous regular Sylvester equation are given in [1, 3, 4].
The singular Sylvester equation over quaternions poses greater challenges. Tian provided solutions for both the homogeneous [5] and inhomogeneous [6] cases, but without explicit proofs, presenting results in forms that obscure the connection to the original problem. Shpakivsky [7] reduced the problem to a system of linear equations over the real number field, while Turner [8] separated real and imaginary components, also transforming the problem into the real domain. Both approaches are algebraically complex, obscuring the link between their structure and the original quaternionic formulation and hindering extension to related problems. Subsequent work, such as that by Janovská and Opfer [1], managed only to characterize the solution space and propose numerical algorithms rather than analytic, closed-form results.
We present both the general and nonzero solutions of the homogeneous singular Sylvester equation in quaternions, establishing the connection to the quaternionic roots. We also restate the solvability conditions with proofs and provide the general solution for the inhomogeneous singular Sylvester equation.
2. Quaternion Fundamentals
[…]
${a}{b}=(\mathrm{Re}\,{a}{})(\mathrm{Re}\,{b}{})-{\mathrm{Im}\,{a}}\!\cdot{\mathrm{Im}\,{b}}+(\mathrm{Re}\,{a}{})({\mathrm{Im}\,{b}})+(\mathrm{Re}\,{b}{})({\mathrm{Im}\,{a}})+{\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}$
and satisfies the following properties
$\mathopen{\lvert}{a}{b}\mathclose{\rvert}=\mathopen{\lvert}{a}\mathclose{\rvert}\kern 1.00006pt\mathopen{\lvert}{b}\mathclose{\rvert},\qquad\mathrm{Re}\,{({a}{b})}=\mathrm{Re}\,{({b}{a})}.$
In particular, if a,b,c{a},{b},{c} are pure quaternions, then it holds
${a}{b}=-{a}\cdot{b}+{a}\times{b},\qquad-{a}^{2}=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}={a}\cdot{a},\qquad{c}\times({a}\times{b})=({c}\cdot{b}){a}-({c}\cdot{a}){b}.$
Lemma 2.1.
For two nonreal quaternions a,b{a},{b} it holds ab=ba{a}{b}={b}{a} if and only if there exists a real number $\lambda$ such that ${\mathrm{Im}\,{a}}=\lambda({\mathrm{Im}\,{b}})$.
Proof.
Since $\mathrm{Re}\,{({a}{b})}=\mathrm{Re}\,{({b}{a})}$ it holds $\mathrm{Re}\,{({a}{b})}-\mathrm{Re}\,{({b}{a})}=0$ such that
${a}{b}-{b}{a}=\mathrm{Im}\,({a}{b})-\mathrm{Im}\,({b}{a})={\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}-{\mathrm{Im}\,{b}}\times{\mathrm{Im}\,{a}}=2\,({\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}),$
which yields the assertion of the lemma. ∎
Lemma 2.2.
For two nonreal quaternions a,b{a},{b} with $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}$ and $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$ it holds ab=ba{a}{b}={b}{a} if and only if either ${\mathrm{Im}\,{a}}={\mathrm{Im}\,{b}}$ or ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$.
Proof.
Given $(\mathrm{Re}\,{a}{})^{2}-(\mathrm{Re}\,{b}{})^{2}=0$ and Lemma 2.1 it holds ab=ba{a}{b}={b}{a} if and only if there exists a real number $\lambda$ such that
$0=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}-\mathopen{\lvert}{b}\mathclose{\rvert}^{2}=({\mathrm{Im}\,{b}})^{2}-({\mathrm{Im}\,{a}})^{2}=({\mathrm{Im}\,{b}})^{2}-(\lambda{\mathrm{Im}\,{b}})^{2}=(1-\lambda^{2})({\mathrm{Im}\,{b}})^{2}.$
Hence, ab=ba{a}{b}={b}{a} holds if and only if $\lambda=\pm 1$, which completes the proof. ∎
Lemma 2.3.
For pure quaternions a,b{a},{b} with $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$ it holds
(2.1) ${a}({a}\times{b})+({a}\times{b}){b}=\big({a}\cdot({b}-{a})\big)({a}+{b}),$
which equals zero if and only if ab=ba{a}{b}={b}{a}.
Proof.
Since ${a}\cdot({a}\times{b})=0$ and ${a}\cdot{a}=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}=\mathopen{\lvert}{b}\mathclose{\rvert}^{2}={b}\cdot{b}$ it holds
${a}({a}\times{b})+({a}\times{b}){b}$ $={a}\times({a}\times{b})-{b}\times({a}\times{b})$
$=({a}\cdot{b}){a}-({a}\cdot{a}){b}-({b}\cdot{b}){a}+({b}\cdot{a}){b}$
$=\big(({a}\cdot{b})-({a}\cdot{a})\big)({a}+{b})$
$=\big({a}\cdot({b}-{a})\big)({a}+{b}),$
Further, by Lemma 2.2 we have ab=ba{a}{b}={b}{a} if and only if either a=b{a}={b} or a=b{a}=-{b} and hence if and only if (2.1) equals zero. ∎
Lemma 2.4.
For two nonreal quaternions a,b{a},{b} it holds ab=ba{a}{b}=-{b}{a} if and only if $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}=0$ and ${\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{b}}=0$.
Proof.
First assume that ab=ba{a}{b}=-{b}{a}. Then we have 0=ab+ba0={a}{b}+{b}{a} and in particular
$0=\mathrm{Im}\,({a}{b})+\mathrm{Im}\,({b}{a})=2[(\mathrm{Re}\,{a}{})({\mathrm{Im}\,{b}})+(\mathrm{Re}\,{b}{})({\mathrm{Im}\,{a}})].$
This implies $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}=0$, since ${a}{b}\neq{b}{a}$ and hence by Lemma 2.1 ${\mathrm{Im}\,{a}}\neq\lambda({\mathrm{Im}\,{b}})$ for all real scalars $\lambda$. Furthermore, we have
$0=\mathrm{Re}\,({a}{b})+\mathrm{Re}\,({b}{a})=2[(\mathrm{Re}\,{a}{})(\mathrm{Re}\,{b}{})-{\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{b}}]=-2({\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{b}}).$
Conversely, assume $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}=0$ and ${\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{b}}=0$. Then
${a}{b}+{b}{a}=\mathrm{Im}\,({a}{b})+\mathrm{Im}\,({b}{a})={\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}+{\mathrm{Im}\,{b}}\times{\mathrm{Im}\,{a}}=0,$
which yields the assertion of the lemma. ∎
Definition 2.5.
Two nonreal quaternions a,b{a},{b} are called similar if there exists a nonzero quaternion p{p} such that ${p}^{-1}{a}{p}={b}$. In that case we write ${a}\sim{b}$.
Remark 2.6.
In Definition 2.5, similarity was introducted only for nonreal quaternions. In general, however, this term may also be defined for real quaternions. For real similar quaternions a,b{a},{b} it holds ${a}=\mathrm{Re}\,\mathopen{(}{a}{p}{p}^{-1}\mathclose{)}=\mathrm{Re}\,\mathopen{(}{p}^{-1}{a}{p}\mathclose{)}={b}$ and further
${p}^{-1}{a}{p}={b}\quad\Longleftrightarrow\quad{a}{p}-{b}{p}=0\quad\Longleftrightarrow\quad({a}-{b}){p}=0,$
such that every nonzero quaternion p{p} possesses this property. This implies, that any quaternion is a solution to the homogeneous singular Sylvester equation, and that the inhomogeneous singular Sylvester equation has no solutions, as will be shown later.
Lemma 2.7.
Two nonreal quaternions a,b{a},{b} are similar if and only if $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}$ and $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$.
Proof.
First assume that ${a}\sim{b}$. Then, there exists a nonzero quaternion p{p} such that $\mathrm{Re}\,{a}{}=\mathrm{Re}\,\mathopen{(}{a}{p}{p}^{-1}\mathclose{)}=\mathrm{Re}\,\mathopen{(}{p}^{-1}{a}{p}\mathclose{)}=\mathrm{Re}\,{b}{}$. Moreover, ap=pb{a}{p}={p}{b} such that
$\mathopen{\lvert}{a}\mathclose{\rvert}\mathopen{\lvert}{p}\mathclose{\rvert}=\mathopen{\lvert}{a}{p}\mathclose{\rvert}=\mathopen{\lvert}{p}{b}\mathclose{\rvert}=\mathopen{\lvert}{p}\mathclose{\rvert}\mathopen{\lvert}{b}\mathclose{\rvert}$
and hence $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$. Conversely, assume that $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}$ and $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$. Then, it holds $\mathrm{Re}\,{a}{}-\mathrm{Re}\,{b}{}=0$ and $\mathopen{\lvert}{b}\mathclose{\rvert}^{2}-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}=0$. Furthermore, we have
(2.2) ${a}{p}-{p}{b}=(\mathrm{Re}\,{a}{}-\mathrm{Re}\,{b}{}){p}+({\mathrm{Im}\,{a}}){p}-{p}({\mathrm{Im}\,{b}})=({\mathrm{Im}\,{a}}){p}-{p}({\mathrm{Im}\,{b}}).$
Now, consider the following cases in order to find a proper nonzero quaternion p{p}. Case 1: Let ${a}{b}\neq{b}{a}$. Then, by Lemma 2.1 it holds ${\mathrm{Im}\,{a}}\neq\lambda({\mathrm{Im}\,{b}})$ for all real numbers $\lambda$. Hence, ${p}={\mathrm{Im}\,{a}}+{\mathrm{Im}\,{b}}$ is nonzero. Case 2: Let ab=ba{a}{b}={b}{a}. Then, by Lemma 2.2 it holds either ${\mathrm{Im}\,{a}}={\mathrm{Im}\,{b}}$ or ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$. Case 2a: Let ${\mathrm{Im}\,{a}}={\mathrm{Im}\,{b}}$. Then, ${p}={\mathrm{Im}\,{a}}+{\mathrm{Im}\,{b}}=2\,{\mathrm{Im}\,{b}}$ is nonzero. Case 2b: Let ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$. Since ${\mathrm{Im}\,{a}}+{\mathrm{Im}\,{b}}=0$, choose an arbitrary nonreal quaternion q{q} with $({\mathrm{Im}\,{a}}){q}\neq{q}({\mathrm{Im}\,{a}})$. Then, ${p}=({\mathrm{Im}\,{a}}){q}+{q}({\mathrm{Im}\,{b}})=({\mathrm{Im}\,{a}}){q}-{q}({\mathrm{Im}\,{a}})$ is nonzero. Given the nonzero quaternions p{p}, for the Cases 1 and 2a we have
(2.3) $({\mathrm{Im}\,{a}}){p}-{p}({\mathrm{Im}\,{b}})=({\mathrm{Im}\,{a}})^{2}-({\mathrm{Im}\,{b}})^{2}=\mathopen{\lvert}{b}\mathclose{\rvert}^{2}-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}=0$
and for Case 2b further
(2.4) $({\mathrm{Im}\,{a}}){p}-{p}({\mathrm{Im}\,{b}})=({\mathrm{Im}\,{a}})^{2}{q}-{q}({\mathrm{Im}\,{b}})^{2}=(\mathopen{\lvert}{b}\mathclose{\rvert}^{2}-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}){q}=0.$
Finally, combining (2.2), (2.3) as well as (2.2), (2.4) yields ap=pb{a}{p}={p}{b} with nonzero p{p}, which completes the proof. ∎
Corollary 2.7.1.
For pure quaternions a,b{a},{b} it holds ${a}\sim{b}$ if and only if $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$ and hence if and only if ${a}\sim-{b}$.
3. Quaternion square roots
When attempting to solve the singular Sylvester equation, constraints resulting from the noncommutativity of the quaternion product inevitably lead to the idea that the solutions are somehow related to quaternion square roots. As will be shown later, these ideas are not unfounded. Square roots of quaternions are defined analogously to square roots of complex numbers and are the solutions to the quadratic equation (3.1) where a{a} represents a nonzero quaternion.
(3.1) ${x}^{2}-{a}=0$
Statements about both the existence and number of such solutions can be found in [9, 10, 11, 12]. According to this, (3.1) has two solutions if and only if either a{a} is a nonreal quaternion or a positive real number. Otherwise, if a{a} is a negative real number, (3.1) has infinitely many solutions. The representation of these solutions will be derived next.
Definition 3.1.
Let a{a} be a nonzero quaternion. The solutions x{x} of (3.1) are called quaternion square root of a{a} and are denoted as $\sqrt{a}$.
Lemma 3.2.
For any nonzero quaternion a{a} (3.1) has the solutions
(3.2) ${x}=\pm\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}\,\frac{{p}}{\mathopen{\lvert}{p}\mathclose{\rvert}}$,
where p{p} is either an arbitrary pure quaternion if a{a} is a negative real number or ${p}={a}+\mathopen{\lvert}{a}\mathclose{\rvert}$ if a{a} is a nonreal quaternion or a positive real number.
Proof.
Given (3.1) we have xx=a{x}{x}={a} such that $\mathopen{\lvert}{x}\mathclose{\rvert}^{2}=\mathopen{\lvert}{a}\mathclose{\rvert}$ and $\mathopen{\lvert}{x}\mathclose{\rvert}=\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}$. Now assume that a{a} is nonreal. Then,
(3.3) $2(\mathrm{Re}\,{x}{}){x}={x}{x}+{x}^{*}{x}={x}{x}+\mathopen{\lvert}{x}\mathclose{\rvert}^{2}={a}+\mathopen{\lvert}{a}\mathclose{\rvert}$.
Assuming $\mathrm{Re}\,{x}{}=0$ we have $0={a}+\mathopen{\lvert}{a}\mathclose{\rvert}$ which contradicts that a{a} is nonreal. Hence, we must have $\mathrm{Re}\,{x}{}\neq 0$ and further
(3.4) $\mathopen{\lvert}2\mathrm{Re}\,{x}{}\mathclose{\rvert}=\frac{\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}}{\mathopen{\lvert}{x}\mathclose{\rvert}}\quad\Longleftrightarrow\quad 2\mathrm{Re}\,{x}{}=\pm\frac{\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}}{\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}}$,
such that substituting (3.4) into (3.3) yields (3.2). Next, assume that a{a} is a real number. Then, from (3.1) we have $0=\mathrm{Im}\,({x}{x})$ and $a=\mathrm{Re}\,({x}{x})$ as well as
$0=\mathrm{Im}\,({x}{x})=2(\mathrm{Re}\,{x}{})({\mathrm{Im}\,{x}})\quad\Longleftrightarrow\quad\mathrm{Re}\,{x}{}=0\text{~or~}{\mathrm{Im}\,{x}}=0.$
Now assume that a{a} is positive. If $\mathrm{Re}\,{x}{}=0$, then $a=\mathrm{Re}\,({x}{x})=-\mathopen{\lvert}{\mathrm{Im}\,{x}}\mathclose{\rvert}^{2}$, which contradicts that a{a} is a positive real number. Hence, we must have ${\mathrm{Im}\,{x}}=0$ and further ${x}=\mathrm{Re}\,{x}{}=\pm\textstyle\sqrt{a}$. Since ${a}=\mathopen{\lvert}{a}\mathclose{\rvert}$ it holds ${a}+\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}$, which yields (3.2). Conversely, assume that a{a} is negative. If ${\mathrm{Im}\,{x}}=0$, then ${a}=\mathrm{Re}\,({x}{x})=\mathopen{\lvert}\mathrm{Re}\,{x}{}\mathclose{\rvert}^{2}$, which contradicts that a{a} is a negative real number. Hence, we must have $\mathrm{Re}\,{x}{}=0$. Further, for any pure quaternion p{p} it holds ${p}{p}=-\mathopen{\lvert}{p}\mathclose{\rvert}^{2}$ such that
${x}{x}=-\mathopen{\lvert}{x}\mathclose{\rvert}^{2}=-\mathopen{\lvert}{a}\mathclose{\rvert}=\frac{{p}{p}}{\mathopen{\lvert}{p}\mathclose{\rvert}^{2}}\mathopen{\lvert}{a}\mathclose{\rvert}$,
which yields (3.2) and completes the proof. ∎
Corollary 3.2.1.
The square roots of any nonzero quaternion a{a} are unit quaternions if and only if quaternion a{a} is unit.
Corollary 3.2.2.
The square roots of any nonreal quaternion a{a} have nonzero real parts, i.e. the square roots of a nonreal quaternion a{a} are not pure.
Corollary 3.2.3.
For any nonreal quaternion a{a} there exist nonzero real numbers $\lambda_{0},\lambda_{1}$ such that
$\sqrt{{a}}=\pm(\lambda_{0}+\lambda_{1}{a}).$
In particular, we have $\lambda_{0}=\lambda_{1}\mathopen{\lvert}{a}\mathclose{\rvert}$ and $\lambda_{1}=\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}/\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}$.
In order to apply Lemma 3.2 to find solutions to the singular Sylvester equation, it is necessary to consider a special case of (3.1). For arbitrary quaternions a,b{a},{b}, for which the product ab{a}{b} is nonzero, the square roots of ab{a}{b} are the solutions to
(3.5) ${x}^{2}-{a}{b}=0.$
If the product ab{a}{b} is nonreal, then (3.5) has exactly two solutions. In addition, if the quaternions a,b{a},{b} have properties that are necessary for the Sylvester equation to be singular, the solutions of (3.5) have a special structure.
Lemma 3.3.
For any quaternions a,b{a},{b} with $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$, for which the quaternion product ab{a}{b} is nonreal, (3.5) has the solutions
(3.6) $x=\pm\frac{{a}({b}+{a}^{*})}{\mathopen{\lvert}{b}+{a}^{*}\mathclose{\rvert}}=\pm\frac{({a}+{b}^{*}){b}}{\mathopen{\lvert}{a}+{b}^{*}\mathclose{\rvert}}$.
Proof.
We have $\mathopen{\lvert}{a}{b}\mathclose{\rvert}=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}={a}{a}^{*}$ such that $\sqrt{\mathopen{\lvert}{a}{b}\mathclose{\rvert}}=\mathopen{\lvert}{a}\mathclose{\rvert}$ and ${a}{b}+\mathopen{\lvert}{a}{b}\mathclose{\rvert}={a}({b}+{a}^{*})$. Then, Lemma 3.2 with ab{a}{b} instead of a{a} yields the left hand solutions of (3.6). The right hand solutions of (3.6) also follow from Lemma 3.2, since $\mathopen{\lvert}{a}{b}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}^{2}={b}^{*}{b}$ such that $\sqrt{\mathopen{\lvert}{a}{b}\mathclose{\rvert}}=\mathopen{\lvert}{b}\mathclose{\rvert}$ and ${a}{b}+\mathopen{\lvert}{a}{b}\mathclose{\rvert}=({a}+{b}^{*}){b}$. ∎
Corollary 3.3.1.
For two quaternions a,b{a},{b} with $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$, for which the quaternion product ab{a}{b} is nonreal, there exist nonzero real numbers $\lambda_{0},\lambda_{1}$ such that
$\sqrt{{a}{b}}=\pm(\lambda_{0}+\lambda_{1}{a}{b}).$
In particular, we have $\lambda_{0}=\lambda_{1}\mathopen{\lvert}{a}\mathclose{\rvert}\mathopen{\lvert}{b}\mathclose{\rvert}$ and $\lambda_{1}=1/\mathopen{\lvert}{a}+{b}^{*}\mathclose{\rvert}=1/\mathopen{\lvert}{b}+{a}^{*}\mathclose{\rvert}$.
4. Solutions to the homogeneous singular Sylvester equation
We consider the general linear function f(x)=axxbf(x)={a}{x}-{x}{b}, which is called singular if there exists a nonzero quaternion x{x} such that $f(x)=0$. If $f$, and consequently the quaternions a,b{a},{b}, satisfy this condition, then
(4.1) ${a}{x}-{x}{b}=0$
is referred to as a homogeneous singular Sylvester equation. In this section, we show the existence of solutions to (4.1) and derive their explicit representation. More precisely, we describe both the general and nontrivial solutions, where in the latter case, the quaternion square root plays a particularly significant role.
Theorem 4.1.
Let a,b{a},{b} be nonreal quaternions. Then, (4.1) has a nonzero solution if and only if ${a}\sim{b}$.
Proof.
By Definition 2.5, there exists a nonzero p{p} such that
${p}^{-1}{a}{p}={b}\quad\Longleftrightarrow\quad{a}{p}={p}{b}\quad\Longleftrightarrow\quad{a}{p}-{p}{b}=0.$
Hence, the nonzero solution x=p{x}={p} yields the assertion of the theorem. ∎
Corollary 4.1.1.
Let a,b{a},{b} be nonreal quaternions. Then, the general linear function f(x)=axxbf(x)={a}{x}-{x}{b} is singular if and only if ${a}\sim{b}$. In that case it holds
(4.2) $f(x)=({\mathrm{Im}\,{a}}){x}-{x}({\mathrm{Im}\,{b}}).$
When singular Sylvester equations are considered, the quaternions a,b{a},{b} may, by property (4.2) and without loss of generality, be considered as pure. We will make use of this assumption in proofs of the following theorems. However, the statements of the theorems will be presented in their general form.
In order to solve (4.1) for similar quaternions a{a} and b{b}, we shall proceed as in the proof of Lemma 2.7. There, we distinguished between the following cases. Case 1: ${a}{b}\neq{b}{a}$ and Case 2: ab=ba{a}{b}={b}{a}. According to Lemma 2.2, the latter holds if and only if either Case 2a: a=b{a}={b} or Case 2b: a=b{a}=-{b}. We show that the solution x=p{x}={p} for Case 1 and Case 2a may be represented in the same way as for Case 2b.
Theorem 4.2.
Let a,b{a},{b} be similar quaternions and q{q} an arbitrary quaternion. Then, the general solution to (4.1) is given as
(4.3) ${x}=({\mathrm{Im}\,{a}}){q}+{q}({\mathrm{Im}\,{b}}).$
Proof.
By Corollary 4.1.1, the quaternions a,b{a},{b} may, without loss of generality, be considered as pure. Let y{y} be an arbitrary solution to (4.1). Then it holds ay=yb{a}{y}={y}{b}. Define q=ay+yb{q}={a}{y}+{y}{b}. Then we have
${x}={a}{q}+{q}{b}={a}^{2}{y}+2{a}{y}{b}+{y}{b}^{2}=\lambda{y},$
where $\lambda=-4\mathopen{\lvert}{a}\mathclose{\rvert}^{2}$. Hence, any solution to (4.1) can be obtained from (4.3) with an appropriate q{q}. Now, let q{q} be an arbitrary quaternion. Then, by substituting (4.3) into (4.1) we obtain
${a}{x}-{x}{b}=\mathopen{\lvert}{b}\mathclose{\rvert}^{2}{q}+{a}{q}{b}-{a}{q}{b}-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}{q}=0.$
Hence, (4.3) is the general solution to (4.1) for any quaternion q{q}. ∎
Theorem 4.3.
Let a,b{a},{b} be similar quaternions and $\lambda,\mu$ arbitrary real numbers with $\mathopen{\lvert}\lambda\mathclose{\rvert}+\mathopen{\lvert}\mu\,\mathrm{Im}\,{({a}+{b})}\mathclose{\rvert}\neq 0$. Then, the nonzero solutions to (4.1) are given as
(4.4) ${x}=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\,\mathrm{Im}\,{({a}+{b})},$
which must satisfy ${\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{x}}=0$ if ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$.
Proof.
By Corollary 4.1.1, the quaternions a,b{a},{b} may, without loss of generality, be considered as pure. We examine the following cases to find the nonzero solutions to (4.1). Case 1: Let ${a}{b}\neq{b}{a}$. Then, by Lemma 2.1, it holds ${\mathrm{Im}\,{a}}\neq\nu({\mathrm{Im}\,{b}})$ for all real numbers $\nu$. Hence, any quaternion q{q} can be given as
(4.5) ${q}=\lambda_{0}+\lambda_{1}{a}+\lambda_{2}{b}+\lambda_{3}({a}\times{b}),$
where $\lambda_{0},\dots,\lambda_{3}$ are arbitrary real numbers. Substituting (4.5) into (4.3) we obtain the general solution to (4.1) as
${x}=\lambda_{0}({a}+{b})-(\lambda_{1}+\lambda_{2})(\mathopen{\lvert}{a}\mathclose{\rvert}^{2}-{a}{b})+\lambda_{3}[{a}({a}\times{b})+({a}\times{b}){b}].$
By applying Lemma 3.3 to $(\mathopen{\lvert}{a}\mathclose{\rvert}^{2}-{a}{b})$, and by Lemma 2.3, we further obtain (4.4). Case 2: Let ab=ba{a}{b}={b}{a}. Then, by Lemma 2.2 either a=b{a}={b} or a=b{a}=-{b} holds. Case 2a: Let a=b{a}={b}. Then, by Theorem 4.2, for any quaternion q{q}, the general solution to (4.1) is given as
${x}={a}{q}+{q}{a}=2{a}\cdot(\mathrm{Re}\,{q}{}-{\mathrm{Im}\,{q}}).$
Let $\lambda,\mu$ be arbitrary real numbers. Then, with ${a}(\mathrm{Re}\,{q}{})=\mu\,({a}+{b})$ and ${a}\cdot{\mathrm{Im}\,{q}}=\lambda\sqrt{{a}{b}^{*}}$, according to Lemma 3.2, we obtain (4.4). Case 2b: Let a=b{a}=-{b}. Then, by Theorem 4.2, and for any quaternion q{q}, the general solution to (4.1) is given as
${x}={a}{q}-{q}{a}=2({a}\times{\mathrm{Im}\,{q}}).$
Hence, x{x} is a pure quaternion that satisfies ${a}\cdot{x}=2{a}\cdot({a}\times{\mathrm{Im}\,{q}})=0$. Moreover, by Lemma 3.2, and with (a+b)=0({a}+{b})=0, we obtain (4.4). Across all cases, the solution in (4.4) is nonzero if and only if $\mathopen{\lvert}\lambda\mathclose{\rvert}+\mathopen{\lvert}\mu({a}+{b})\mathclose{\rvert}\neq 0$, which completes the proof. ∎
5. Solutions to the inhomogeneous singular Sylvester equation
In this section we present the conditions for solvability and the general solution to the inhomogeneous singular Sylvester equation. For similar quaternions a,b{a},{b} and a nonzero quaternion c{c}, the inhomogeneous singular Sylvester equation is given as
(5.1) ${a}{x}-{x}{b}={c}.$
If a,b{a},{b} are real numbers, (5.1) has no solution, since (ab)x=0({a}-{b}){x}=0 for all quaternions x{x}. Therefore, in the following theorems, we assume that a,b{a},{b} are nonreal quaternions. Analogous to Lemma 2.7, we investigate the solutions to (5.1) for the following casses. Case 1: ${a}{b}\neq{b}{a}$, Case 2: ab=ba{a}{b}={b}{a}. Then, by Lemma 2.1 it holds either Case 2a: a=b{a}={b} or Case 2b: ${a}={b}^{*}$. The solvability condition influences the structure of the solution.
Theorem 5.1.
Let a,b,c{a},{b},{c} be nonzero quaternions with ${a}\sim{b}$. Then, (5.1) has a solution if and only if ${a}{c}={c}{b}^{*}$.
Proof.
By Corollary 4.1.1, we may assume that a,b{a},{b} are pure quaternions. Since $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$, we have
${a}{c}={a}({a}{x}-{x}{b})=-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}{x}-{a}{x}{b}=-{a}{x}{b}-{x}\mathopen{\lvert}{b}\mathclose{\rvert}^{2}=-({a}{x}-{x}{b}){b}=-{c}{b},$
and hence the assertion of the theorem. ∎
Corollary 5.1.1.
Case 2a: Let a=b{a}={b}. Then, (5.1) has a solution if and only if $({\mathrm{Im}\,{a}}){c}=-{c}({\mathrm{Im}\,{a}})$. Case 2b: Let ${a}={b}^{*}$. Then, (5.1) has a solution if and only if ac=ca{a}{c}={c}{a}.
Theorem 5.2.
Let a,b,c{a},{b},{c} be nonzero quaternions with ${a}\sim{b}$ and ${a}{c}={c}{b}^{*}$. Then, for any quaternion q{q}, the general solution to (5.1) is given as
(5.2) ${x}=({\mathrm{Im}\,{a}})\frac{{q}-{c}}{4\mathopen{\lvert}{\mathrm{Im}\,{a}}\mathclose{\rvert}^{2}}+\frac{{q}+{c}}{4\mathopen{\lvert}{\mathrm{Im}\,{b}}\mathclose{\rvert}^{2}}({\mathrm{Im}\,{b}}).$
Proof.
By Corollary 4.1.1, we may assume that a,b{a},{b} are pure quaternions. Substituting ${c}={a}{a}^{-1}{c}$ and ${c}={c}{b}^{-1}{b}$ into (5.1) yields
(5.3) $0={a}{x}-{x}{b}-\frac{1}{2}({a}{a}^{-1}{c}+{c}{b}^{-1}{b})={a}({x}+\frac{{a}{c}}{2\mathopen{\lvert}{a}\mathclose{\rvert}^{2}})-({x}-\frac{{c}{b}}{2\mathopen{\lvert}{b}\mathclose{\rvert}^{2}}){b}.$
Applying Theorem 4.2 to the homogeneous singular Sylvester equation (5.3), we have
(5.4) ${x}+\frac{{a}{c}}{2\mathopen{\lvert}{a}\mathclose{\rvert}^{2}}={a}{q}+{q}{b}\qquad\Longleftrightarrow\qquad{x}={a}{q}+{q}{b}-\frac{{a}{c}}{2\mathopen{\lvert}{a}\mathclose{\rvert}^{2}},$
for any quaternion q{q}. Substituting ${a}{c}=\frac{1}{2}({a}{c}-{c}{b})$ into (5.4) yields (5.2). ∎
Corollary 5.2.1.
Let a,b,c{a},{b},{c} be nonzero quaternions with ${a}\sim{b}$ and ${a}{c}={c}{b}^{*}$. Then, for arbitrary real numbers $\lambda,\mu$ the general solution to (5.1) is given as
${x}=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\mathrm{Im}\,({a}+{b})+\frac{{c}}{4\mathopen{\lvert}{\mathrm{Im}\,{a}}\mathclose{\rvert}^{2}}\mathrm{Im}\,({b}-{a}).$
6. Conclusion
We have presented closed-form solutions to the homogeneous and inhomogeneous singular Sylvester equations in quaternions, together with a restatement of the corresponding solvability conditions. In addition, we derived an explicit formula for quaternion roots and showed how the closed-form solution naturally incorporates the quaternion square root. This structural connection provides a pathway for extending the methodology to a broader class of problems in quaternion analysis.
Acknowledgement
The authors would like to thank the Federal Ministry of Research, Technology, and Space (BMFTR) for its support as part of the research program Communication Systems “Souverän. Digital. Vernetzt.”. Joint project 6G-life, project identification number: 16KIS2413K and the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/2 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of TUD Dresden University of Technology.