### 27 Mar 2026 v4

# On solutions of singular Sylvester equations in quaternions

Abstract.
The quaternionic equations axxb=0{a}{x}-{x}{b}=0 and axxb=c{a}{x}-{x}{b}=c are investigated, which are called homogeneous and inhomogeneous Sylvester equations, respectively. Conditions for the existence of solutions are provided. In addition, the general and nonzero solutions to these equations are derived applying quaternion square roots.
1. Introduction
A general linear function of a quaternion variable x{x} has the form [1]
$f_{n}({x})={a}_{1}{x}{b}_{1}+{a}_{2}{x}{b}_{2}+\dots+{a}_{n}{x}{b}_{n},$
where the coefficients ${a}_{j},{b}_{j}$ for $j=1,2,\dots,n$ are nonzero quaternions. Due to the noncommutativity of the quaternion product, $f_{n}$ is in general not homogeneous and hence not linear. However, $f_{n}$ is linear over the real numbers [1]. For $n=2$ and without loss of generality $f_{n}$ may be rewritten as
$f({x})={a}{x}-{x}{b}.$
The function $f$ is called singular if there exists a nonzero quaternion x{x} for which f(x)=0f({x})=0, otherwise it is called regular. Given a quaternion c{c} the equation
$f({x})={a}{x}-{x}{b}={c},$
is referred to as the Sylvester equation. The equation is called homogeneous if c=0{c}=0 and inhomogeneous otherwise. The Sylvester equation is called singular if the associated function $f$ is singular, otherwise regular.
Johnson [2] distinguished between the regular and singular cases and provided topological proofs for the existence and uniqueness of solutions over division rings, which directly apply to quaternions. A regular Sylvester equation has a unique solution for every quaternion c{c}. Hence, the homogeneous regular Sylvester equation has only the trivial solution x=0{x}=0. Closed-form expressions for the inhomogeneous regular Sylvester equation are given in [1, 3, 4].
The singular Sylvester equation over quaternions poses greater challenges. Rodman [5] provided a comprehensive matrix-theoretic foundation for quaternionic linear algebra, including detailed analyses of eigenvalues and similarity transformations, yet stopped short of providing explicit closed-form solutions for the singular Sylvester equation. Tian provided solutions for both the homogeneous [6] and inhomogeneous [7] cases, but without explicit proofs, presenting results in forms that obscure the connection to the original problem. Shpakivsky [8] reduced the problem to a system of linear equations over the real number field, while Turner [9] separated real and imaginary components, also transforming the problem into the real domain. Both approaches are algebraically complex, obscuring the link between their structure and the original quaternionic formulation and hindering extension to related problems. Subsequent work, such as that by Janovská and Opfer [1], managed only to characterize the solution space and propose numerical algorithms rather than analytic, closed-form results.
The main contribution of this work is the derivation of the analytic, closed-form nonzero solutions for the homogeneous and the general solution of the inhomogeneous singular Sylvester equations, as presented in Theorems 4.3 and 5.3, respectively. By establishing a structural connection to quaternionic square roots, we uncover a direct geometric decomposition of the solution space. This formulation facilitates a deeper understanding of the solution’s orientation relative to the axes of a{a} and b{b}, which is essential for engineering applications involving spatial rotations.
2. Quaternion Fundamentals
[…]
${a}{b}=(\mathrm{Re}\,{a}{})(\mathrm{Re}\,{b}{})-{\mathrm{Im}\,{a}}\!\cdot{\mathrm{Im}\,{b}}+(\mathrm{Re}\,{a}{})({\mathrm{Im}\,{b}})+(\mathrm{Re}\,{b}{})({\mathrm{Im}\,{a}})+{\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}$
where the symbols (,×)($\cdot\,$$\times$) represent the vector scalar and cross product, respectively. The quaternion product satisfies the following properties
$\mathopen{\lvert}{a}{b}\mathclose{\rvert}=\mathopen{\lvert}{a}\mathclose{\rvert}\kern 1.00006pt\mathopen{\lvert}{b}\mathclose{\rvert},\qquad\mathrm{Re}\,{({a}{b})}=\mathrm{Re}\,{({b}{a})}.$
In particular, if a,b,c{a},{b},{c} are pure quaternions, then it holds
${a}{b}=-{a}\cdot{b}+{a}\times{b},\qquad-{a}^{2}=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}={a}\cdot{a},\qquad{c}\times({a}\times{b})=({c}\cdot{b}){a}-({c}\cdot{a}){b}.$
Lemma 2.1.
For two nonreal quaternions a,b{a},{b} it holds ab=ba{a}{b}={b}{a} if and only if there exists a real number $\lambda$ such that ${\mathrm{Im}\,{a}}=\lambda({\mathrm{Im}\,{b}})$.
Proof.
Since $\mathrm{Re}\,{({a}{b})}=\mathrm{Re}\,{({b}{a})}$ it holds $\mathrm{Re}\,{({a}{b})}-\mathrm{Re}\,{({b}{a})}=0$ such that
${a}{b}-{b}{a}=\mathrm{Im}\,({a}{b})-\mathrm{Im}\,({b}{a})={\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}-{\mathrm{Im}\,{b}}\times{\mathrm{Im}\,{a}}=2\,({\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}),$
which yields the assertion of the lemma. ∎
Lemma 2.2.
For two nonreal quaternions a,b{a},{b} with $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}$ and $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$ it holds ab=ba{a}{b}={b}{a} if and only if either ${\mathrm{Im}\,{a}}={\mathrm{Im}\,{b}}$ or ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$.
Proof.
Given $(\mathrm{Re}\,{a}{})^{2}-(\mathrm{Re}\,{b}{})^{2}=0$ and Lemma 2.1 it holds ab=ba{a}{b}={b}{a} if and only if there exists a real number $\lambda$ such that
$0=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}-\mathopen{\lvert}{b}\mathclose{\rvert}^{2}=({\mathrm{Im}\,{b}})^{2}-({\mathrm{Im}\,{a}})^{2}=({\mathrm{Im}\,{b}})^{2}-(\lambda{\mathrm{Im}\,{b}})^{2}=(1-\lambda^{2})({\mathrm{Im}\,{b}})^{2}.$
Hence, ab=ba{a}{b}={b}{a} holds if and only if $\lambda=\pm 1$, which completes the proof. ∎
Lemma 2.3.
For pure quaternions a,b{a},{b} with $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$ it holds
(2.1) ${a}({a}\times{b})+({a}\times{b}){b}=\big({a}\cdot({b}-{a})\big)({a}+{b}),$
which equals zero if and only if ab=ba{a}{b}={b}{a}.
Proof.
Since ${a}\cdot({a}\times{b})=0$ and ${a}\cdot{a}=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}=\mathopen{\lvert}{b}\mathclose{\rvert}^{2}={b}\cdot{b}$ it holds
${a}({a}\times{b})+({a}\times{b}){b}$ $={a}\times({a}\times{b})-{b}\times({a}\times{b})$
$=({a}\cdot{b}){a}-({a}\cdot{a}){b}-({b}\cdot{b}){a}+({b}\cdot{a}){b}$
$=\big(({a}\cdot{b})-({a}\cdot{a})\big)({a}+{b})$
$=\big({a}\cdot({b}-{a})\big)({a}+{b}),$
Further, by Lemma 2.2 we have ab=ba{a}{b}={b}{a} if and only if either a=b{a}={b} or a=b{a}=-{b} and hence if and only if (2.1) equals zero. ∎
Lemma 2.4.
For two nonreal quaternions a,b{a},{b} it holds ab=ba{a}{b}=-{b}{a} if and only if $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}=0$ and ${\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{b}}=0$.
Proof.
First assume that ab=ba{a}{b}=-{b}{a}. Then we have 0=ab+ba0={a}{b}+{b}{a} and in particular
$0=\mathrm{Im}\,({a}{b})+\mathrm{Im}\,({b}{a})=2[(\mathrm{Re}\,{a}{})({\mathrm{Im}\,{b}})+(\mathrm{Re}\,{b}{})({\mathrm{Im}\,{a}})].$
This implies $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}=0$, since ${a}{b}\neq{b}{a}$ and hence by Lemma 2.1 ${\mathrm{Im}\,{a}}\neq\lambda({\mathrm{Im}\,{b}})$ for all real scalars $\lambda$. Furthermore, we have
$0=\mathrm{Re}\,({a}{b})+\mathrm{Re}\,({b}{a})=2[(\mathrm{Re}\,{a}{})(\mathrm{Re}\,{b}{})-{\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{b}}]=-2({\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{b}}).$
Conversely, assume $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}=0$ and ${\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{b}}=0$. Then
${a}{b}+{b}{a}=\mathrm{Im}\,({a}{b})+\mathrm{Im}\,({b}{a})={\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}+{\mathrm{Im}\,{b}}\times{\mathrm{Im}\,{a}}=0,$
which yields the assertion of the lemma. ∎
Definition 2.5.
Two nonreal quaternions a,b{a},{b} are called similar if there exists a nonzero quaternion p{p} such that ${p}^{-1}{a}{p}={b}$. In that case we write ${a}\sim{b}$.
Remark 2.6.
In Definition 2.5, similarity was restricted to nonreal quaternions. However, the concept extends naturally to the real case. For real similar quaternions a,b{a},{b} it holds ${a}=\mathrm{Re}\,\mathopen{(}{a}{p}{p}^{-1}\mathclose{)}=\mathrm{Re}\,\mathopen{(}{p}^{-1}{a}{p}\mathclose{)}={b}$ and further
${p}^{-1}{a}{p}={b}\quad\Longleftrightarrow\quad{a}{p}-{b}{p}=0\quad\Longleftrightarrow\quad({a}-{b}){p}=0,$
such that every nonzero quaternion p{p} possesses this property. This implies that in the real case, every quaternion is a solution to the homogeneous singular Sylvester equation. Conversely, the inhomogeneous singular Sylvester equation has no solutions, as will be shown later.
Lemma 2.7.
Two nonreal quaternions a,b{a},{b} are similar if and only if $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}$ and $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$.
Proof.
First assume that ${a}\sim{b}$. Then, there exists a nonzero quaternion p{p} such that $\mathrm{Re}\,{a}{}=\mathrm{Re}\,\mathopen{(}{a}{p}{p}^{-1}\mathclose{)}=\mathrm{Re}\,\mathopen{(}{p}^{-1}{a}{p}\mathclose{)}=\mathrm{Re}\,{b}{}$. Moreover, ap=pb{a}{p}={p}{b} such that
$\mathopen{\lvert}{a}\mathclose{\rvert}\mathopen{\lvert}{p}\mathclose{\rvert}=\mathopen{\lvert}{a}{p}\mathclose{\rvert}=\mathopen{\lvert}{p}{b}\mathclose{\rvert}=\mathopen{\lvert}{p}\mathclose{\rvert}\mathopen{\lvert}{b}\mathclose{\rvert}$
and hence $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$. Conversely, assume that $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}$ and $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$. Then, it holds $\mathrm{Re}\,{a}{}-\mathrm{Re}\,{b}{}=0$ and $\mathopen{\lvert}{b}\mathclose{\rvert}^{2}-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}=0$. Furthermore, we have
(2.2) ${a}{p}-{p}{b}=(\mathrm{Re}\,{a}{}-\mathrm{Re}\,{b}{}){p}+({\mathrm{Im}\,{a}}){p}-{p}({\mathrm{Im}\,{b}})=({\mathrm{Im}\,{a}}){p}-{p}({\mathrm{Im}\,{b}}).$
Now, consider the following cases in order to find a proper nonzero quaternion p{p}. Case 1: Let ${a}{b}\neq{b}{a}$. Then, by Lemma 2.1 it holds ${\mathrm{Im}\,{a}}\neq\lambda({\mathrm{Im}\,{b}})$ for all real numbers $\lambda$. Hence, ${p}={\mathrm{Im}\,{a}}+{\mathrm{Im}\,{b}}$ is nonzero. Case 2: Let ab=ba{a}{b}={b}{a}. Then, by Lemma 2.2 it holds either ${\mathrm{Im}\,{a}}={\mathrm{Im}\,{b}}$ or ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$. Case 2a: Let ${\mathrm{Im}\,{a}}={\mathrm{Im}\,{b}}$. Then, ${p}={\mathrm{Im}\,{a}}+{\mathrm{Im}\,{b}}=2\,{\mathrm{Im}\,{b}}$ is nonzero. Case 2b: Let ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$. Since ${\mathrm{Im}\,{a}}+{\mathrm{Im}\,{b}}=0$, choose an arbitrary nonreal quaternion q{q} with $({\mathrm{Im}\,{a}}){q}\neq{q}({\mathrm{Im}\,{a}})$. Then, ${p}=({\mathrm{Im}\,{a}}){q}+{q}({\mathrm{Im}\,{b}})=({\mathrm{Im}\,{a}}){q}-{q}({\mathrm{Im}\,{a}})$ is nonzero. Given the nonzero quaternions p{p}, for the Cases 1 and 2a we have
(2.3) $({\mathrm{Im}\,{a}}){p}-{p}({\mathrm{Im}\,{b}})=({\mathrm{Im}\,{a}})^{2}-({\mathrm{Im}\,{b}})^{2}=\mathopen{\lvert}{b}\mathclose{\rvert}^{2}-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}=0$
and for Case 2b further
(2.4) $({\mathrm{Im}\,{a}}){p}-{p}({\mathrm{Im}\,{b}})=({\mathrm{Im}\,{a}})^{2}{q}-{q}({\mathrm{Im}\,{b}})^{2}=(\mathopen{\lvert}{b}\mathclose{\rvert}^{2}-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}){q}=0.$
Finally, combining (2.2), (2.3) as well as (2.2), (2.4) yields ap=pb{a}{p}={p}{b} with nonzero p{p}, which completes the proof. ∎
Corollary 2.7.1.
For pure quaternions a,b{a},{b} it holds ${a}\sim{b}$ if and only if $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$ and hence if and only if ${a}\sim-{b}$.
Remark 2.8.
The result concerning the similarity of the quaternions a{a} and b{b}, as stated in Lemma 2.7, is already present in the equivalence between (1) and (5) of Theorem 2.2.6 in [5]. However, the proof of this theorem does not provide a formula for explicitly determining a quaternion p{p} that establishes the similarity between a{a} and b{b}. In contrast, the proof of Lemma 2.7 explicitly constructs such a p{p} by considering cases 1, 2a, and 2b. This distinction is crucial as it enables the explicit determination of the nonzero solutions to the homogeneous singular Sylvester equation in Theorem 4.3.
3. Quaternion square roots
When attempting to solve the singular Sylvester equation, constraints resulting from the noncommutativity of the quaternion product inevitably lead to the idea that the solutions are somehow related to quaternion square roots. As will be shown later, these ideas are not unfounded. Square roots of quaternions are defined analogously to square roots of complex numbers and are the solutions to the quadratic equation (3.1) where a{a} represents a nonzero quaternion.
(3.1) ${x}^{2}-{a}=0$
Statements about both the existence and number of such solutions can be found in [10, 11, 12, 13]. According to this, (3.1) has two solutions if and only if either a{a} is a nonreal quaternion or a positive real number. Otherwise, if a{a} is a negative real number, (3.1) has infinitely many solutions. The representation of these solutions will be derived next.
Definition 3.1.
Let a{a} be a nonzero quaternion. The solutions x{x} of (3.1) are called quaternion square root of a{a} and are denoted as $\sqrt{a}$.
Lemma 3.2.
For a nonzero quaternion a{a} (3.1) has the solutions
(3.2) ${x}=\pm\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}\,\frac{{p}}{\mathopen{\lvert}{p}\mathclose{\rvert}}$,
where p{p} is either an arbitrary pure quaternion if a{a} is a negative real number or ${p}={a}+\mathopen{\lvert}{a}\mathclose{\rvert}$ if a{a} is a nonreal quaternion or a positive real number.
Proof.
Given (3.1) we have xx=a{x}{x}={a} such that $\mathopen{\lvert}{x}\mathclose{\rvert}^{2}=\mathopen{\lvert}{a}\mathclose{\rvert}$ and $\mathopen{\lvert}{x}\mathclose{\rvert}=\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}$. Now assume that a{a} is nonreal. Then,
(3.3) $2(\mathrm{Re}\,{x}{}){x}={x}{x}+{x}^{*}{x}={x}{x}+\mathopen{\lvert}{x}\mathclose{\rvert}^{2}={a}+\mathopen{\lvert}{a}\mathclose{\rvert}$.
Assuming $\mathrm{Re}\,{x}{}=0$ we have $0={a}+\mathopen{\lvert}{a}\mathclose{\rvert}$ which contradicts that a{a} is nonreal. Hence, we must have $\mathrm{Re}\,{x}{}\neq 0$ and further
(3.4) $\mathopen{\lvert}2\mathrm{Re}\,{x}{}\mathclose{\rvert}=\frac{\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}}{\mathopen{\lvert}{x}\mathclose{\rvert}}\quad\Longleftrightarrow\quad 2\mathrm{Re}\,{x}{}=\pm\frac{\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}}{\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}}$,
such that substituting (3.4) into (3.3) yields (3.2). Next, assume that a{a} is a real number. Then, from (3.1) we have $0=\mathrm{Im}\,({x}{x})$ and $a=\mathrm{Re}\,({x}{x})$ as well as
$0=\mathrm{Im}\,({x}{x})=2(\mathrm{Re}\,{x}{})({\mathrm{Im}\,{x}})\quad\Longleftrightarrow\quad\mathrm{Re}\,{x}{}=0\text{~or~}{\mathrm{Im}\,{x}}=0.$
Now assume that a{a} is positive. If $\mathrm{Re}\,{x}{}=0$, then $a=\mathrm{Re}\,({x}{x})=-\mathopen{\lvert}{\mathrm{Im}\,{x}}\mathclose{\rvert}^{2}$, which contradicts that a{a} is a positive real number. Hence, we must have ${\mathrm{Im}\,{x}}=0$ and further ${x}=\mathrm{Re}\,{x}{}=\pm\textstyle\sqrt{a}$. Since ${a}=\mathopen{\lvert}{a}\mathclose{\rvert}$ it holds ${a}+\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}$, which yields (3.2). Conversely, assume that a{a} is negative. If ${\mathrm{Im}\,{x}}=0$, then ${a}=\mathrm{Re}\,({x}{x})=\mathopen{\lvert}\mathrm{Re}\,{x}{}\mathclose{\rvert}^{2}$, which contradicts that a{a} is a negative real number. Hence, we must have $\mathrm{Re}\,{x}{}=0$. Further, for every pure quaternion p{p} it holds ${p}{p}=-\mathopen{\lvert}{p}\mathclose{\rvert}^{2}$ such that
${x}{x}=-\mathopen{\lvert}{x}\mathclose{\rvert}^{2}=-\mathopen{\lvert}{a}\mathclose{\rvert}=\frac{{p}{p}}{\mathopen{\lvert}{p}\mathclose{\rvert}^{2}}\mathopen{\lvert}{a}\mathclose{\rvert}$,
which yields (3.2) and completes the proof. ∎
Corollary 3.2.1.
The square roots of a nonzero quaternion a{a} are unit quaternions if and only if quaternion a{a} is unit.
Corollary 3.2.2.
The square roots of a nonreal quaternion a{a} have nonzero real parts, i.e. the square roots of a nonreal quaternion a{a} are not pure.
Corollary 3.2.3.
For a nonreal quaternion a{a} there exist nonzero real numbers $\lambda_{0},\lambda_{1}$ such that
$\sqrt{{a}}=\pm(\lambda_{0}+\lambda_{1}{a}).$
In particular, we have $\lambda_{0}=\lambda_{1}\mathopen{\lvert}{a}\mathclose{\rvert}$ and $\lambda_{1}=\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}/\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}$.
Remark 3.3.
A formula for calculating the quaternion square root is already provided in a similar form in Theorem 2.5.2 of [5]. However, the representation used in Lemma 3.2 is more suitable for practical applications in engineering systems, as the general form according to (3.2) applies to arbitrary quaternions a{a}. The only difference lies in the specific representation of quaternion q{q}, which depends on a{a}. In [5] however, completely different representations depending on a{a} were provided.
In order to apply Lemma 3.2 to find solutions to the singular Sylvester equation, it is necessary to consider a special case of (3.1). For arbitrary quaternions a,b{a},{b}, for which the product ab{a}{b} is nonzero, the square roots of ab{a}{b} are the solutions to
(3.5) ${x}^{2}-{a}{b}=0.$
If the product ab{a}{b} is nonreal, then (3.5) has exactly two solutions. In addition, if the quaternions a,b{a},{b} have properties that are necessary for the Sylvester equation to be singular, the solutions of (3.5) have a special structure.
Lemma 3.4.
For arbitrary quaternions a,b{a},{b} with $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$, for which the quaternion product ab{a}{b} is nonreal, (3.5) has the solutions
(3.6) $x=\pm\frac{{a}({b}+{a}^{*})}{\mathopen{\lvert}{b}+{a}^{*}\mathclose{\rvert}}=\pm\frac{({a}+{b}^{*}){b}}{\mathopen{\lvert}{a}+{b}^{*}\mathclose{\rvert}}$.
Proof.
We have $\mathopen{\lvert}{a}{b}\mathclose{\rvert}=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}={a}{a}^{*}$ such that $\sqrt{\mathopen{\lvert}{a}{b}\mathclose{\rvert}}=\mathopen{\lvert}{a}\mathclose{\rvert}$ and ${a}{b}+\mathopen{\lvert}{a}{b}\mathclose{\rvert}={a}({b}+{a}^{*})$. Then, Lemma 3.2 with ab{a}{b} instead of a{a} yields the left hand solutions of (3.6). The right hand solutions of (3.6) also follow from Lemma 3.2, since $\mathopen{\lvert}{a}{b}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}^{2}={b}^{*}{b}$ such that $\sqrt{\mathopen{\lvert}{a}{b}\mathclose{\rvert}}=\mathopen{\lvert}{b}\mathclose{\rvert}$ and ${a}{b}+\mathopen{\lvert}{a}{b}\mathclose{\rvert}=({a}+{b}^{*}){b}$. ∎
Corollary 3.4.1.
For two quaternions a,b{a},{b} with $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$, for which the quaternion product ab{a}{b} is nonreal, there exist nonzero real numbers $\lambda_{0},\lambda_{1}$ such that
$\sqrt{{a}{b}}=\pm(\lambda_{0}+\lambda_{1}{a}{b}).$
In particular, we have $\lambda_{0}=\lambda_{1}\mathopen{\lvert}{a}\mathclose{\rvert}\mathopen{\lvert}{b}\mathclose{\rvert}$ and $\lambda_{1}=1/\mathopen{\lvert}{a}+{b}^{*}\mathclose{\rvert}=1/\mathopen{\lvert}{b}+{a}^{*}\mathclose{\rvert}$.
4. Solutions to the homogeneous singular Sylvester equation
We consider the general linear function f(x)=axxbf(x)={a}{x}-{x}{b}, which is called singular if there exists a nonzero quaternion x{x} such that $f(x)=0$. If $f$, and consequently the quaternions a,b{a},{b}, satisfy this condition, then
(4.1) ${a}{x}-{x}{b}=0$
is referred to as a homogeneous singular Sylvester equation. In this section, we restate the existance and the general solution to (4.1) previously presented in [6] and provide constructive proofs. Moving beyond existing theory, we also provide a novel characterization of the nonzero solutions to (4.1) by applying the quaternion square root, presented in Theorem 4.3.
Lemma 4.1.
Let a,b{a},{b} be nonreal quaternions. Then, (4.1) has a nonzero solution if and only if ${a}\sim{b}$.
Proof.
By Definition 2.5, there exists a nonzero p{p} such that
${p}^{-1}{a}{p}={b}\quad\Longleftrightarrow\quad{a}{p}={p}{b}\quad\Longleftrightarrow\quad{a}{p}-{p}{b}=0.$
Hence, the nonzero solution x=p{x}={p} yields the assertion of the theorem. ∎
Corollary 4.1.1.
Let a,b{a},{b} be nonreal quaternions. Then, the general linear function f(x)=axxbf(x)={a}{x}-{x}{b} is singular if and only if ${a}\sim{b}$. In that case it holds
(4.2) $f(x)=({\mathrm{Im}\,{a}}){x}-{x}({\mathrm{Im}\,{b}}).$
When singular Sylvester equations are considered, the quaternions a,b{a},{b} may, by property (4.2) and without loss of generality, be considered as pure. We will make use of this assumption in proofs of the following theorems. However, the statements of the theorems will be presented in their general form.
In order to solve (4.1) for similar quaternions a{a} and b{b}, we shall proceed as in the proof of Lemma 2.7. There, we distinguished between the following cases. Case 1: ${a}{b}\neq{b}{a}$ and Case 2: ab=ba{a}{b}={b}{a}. According to Lemma 2.2, the latter holds if and only if either Case 2a: a=b{a}={b} or Case 2b: a=b{a}=-{b}. We show that the solution x=p{x}={p} for Case 1 and Case 2a may be represented in the same way as for Case 2b.
Lemma 4.2.
Let a,b{a},{b} be similar quaternions and q{q} an arbitrary quaternion. Then, the general solution to (4.1) is given as
(4.3) ${x}=({\mathrm{Im}\,{a}}){q}+{q}({\mathrm{Im}\,{b}}).$
Proof.
By Corollary 4.1.1, the quaternions a,b{a},{b} may, without loss of generality, be considered as pure. Let y{y} be an arbitrary solution to (4.1). Then it holds ay=yb{a}{y}={y}{b}. Define q=ay+yb{q}={a}{y}+{y}{b}. Then we have
${x}={a}{q}+{q}{b}={a}^{2}{y}+2{a}{y}{b}+{y}{b}^{2}=\lambda{y},$
where $\lambda=-4\mathopen{\lvert}{a}\mathclose{\rvert}^{2}$. Hence, every solution to (4.1) can be obtained from (4.3) with an appropriate q{q}. Now, let q{q} be an arbitrary quaternion. Then, by substituting (4.3) into (4.1), we obtain
${a}{x}-{x}{b}=\mathopen{\lvert}{b}\mathclose{\rvert}^{2}{q}+{a}{q}{b}-{a}{q}{b}-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}{q}=0.$
Hence, (4.3) is the general solution to (4.1) for an arbitrary quaternion q{q}. ∎
Theorem 4.3.
Let a,b{a},{b} be similar quaternions and $\lambda,\mu$ arbitrary real numbers with $\mathopen{\lvert}\lambda\mathclose{\rvert}+\mathopen{\lvert}\mu\,\mathrm{Im}\,{({a}+{b})}\mathclose{\rvert}\neq 0$. Then, the nonzero solutions to (4.1) are given as
(4.4) ${x}=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\,\mathrm{Im}\,{({a}+{b})},$
which must satisfy ${\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{x}}=0$ if ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$.
Proof.
By Corollary 4.1.1, the quaternions a,b{a},{b} may, without loss of generality, be considered as pure. We examine the following cases to find the nonzero solutions to (4.1). Case 1: Let ${a}{b}\neq{b}{a}$. Then, by Lemma 2.1, it holds ${\mathrm{Im}\,{a}}\neq\nu({\mathrm{Im}\,{b}})$ for all real numbers $\nu$. Hence, every quaternion q{q} can be given as
(4.5) ${q}=\lambda_{0}+\lambda_{1}{a}+\lambda_{2}{b}+\lambda_{3}({a}\times{b}),$
where $\lambda_{0},\dots,\lambda_{3}$ are arbitrary real numbers. Substituting (4.5) into (4.3) we obtain the general solution to (4.1) as
${x}=\lambda_{0}({a}+{b})-(\lambda_{1}+\lambda_{2})(\mathopen{\lvert}{a}\mathclose{\rvert}^{2}-{a}{b})+\lambda_{3}[{a}({a}\times{b})+({a}\times{b}){b}].$
By applying Lemma 3.4 to $(\mathopen{\lvert}{a}\mathclose{\rvert}^{2}-{a}{b})$, and by Lemma 2.3, we further obtain (4.4). Case 2: Let ab=ba{a}{b}={b}{a}. Then, by Lemma 2.2 either a=b{a}={b} or a=b{a}=-{b} holds. Case 2a: Let a=b{a}={b}. Then, by Lemma 4.2, for an arbitrary quaternion q{q}, the general solution to (4.1) is given as
${x}={a}{q}+{q}{a}=2{a}(\mathrm{Re}\,{q}{})-2{a}\cdot{\mathrm{Im}\,{q}}.$
Then, with real numbers $\mu=\mathrm{Re}\,{q}{}$ and $\lambda=-2({a}\cdot{\mathrm{Im}\,{q}})/\mathopen{\lvert}{a}\mathclose{\rvert}$ it follows
$2{a}(\mathrm{Re}\,{q}{})=\mu\,({a}+{b}),\qquad-2{a}\cdot{\mathrm{Im}\,{q}}=\lambda\mathopen{\lvert}{a}\mathclose{\rvert}=\lambda\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}^{2}}=\lambda\sqrt{{a}{b}^{*}},$
which yields (4.4). Case 2b: Let a=b{a}=-{b}. Then, by Lemma 4.2, for an arbitrary quaternion q{q}, the general solution to (4.1) is given as
${x}={a}{q}-{q}{a}=2({a}\times{\mathrm{Im}\,{q}}).$
Since x{x} is a pure quaternion that satisfies ${a}\cdot{x}=2{a}\cdot({a}\times{\mathrm{Im}\,{q}})=0$, with $\lambda=\mathopen{\lvert}{x}\mathclose{\rvert}/\mathopen{\lvert}{a}\mathclose{\rvert}$ and by Lemma 3.2 we have
${x}=\lambda\mathopen{\lvert}{a}\mathclose{\rvert}\frac{{x}}{\mathopen{\lvert}{x}\mathclose{\rvert}}=\lambda\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}^{2}}\frac{{x}}{\mathopen{\lvert}{x}\mathclose{\rvert}}=\lambda\sqrt{-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}}=\lambda\sqrt{{a}{b}^{*}},$
where the expression on the right hand side must satisfy ${a}\cdot\sqrt{{a}{b}^{*}}=0$. Then, with (a+b)=0({a}+{b})=0 we obtain (4.4). Across all cases, the solution in (4.4) is nonzero if and only if $\mathopen{\lvert}\lambda\mathclose{\rvert}+\mathopen{\lvert}\mu({a}+{b})\mathclose{\rvert}\neq 0$, which completes the proof. ∎
Remark 4.4.
The representation in (4.4) identifies the quaternionic directions that span the solution space of (4.1). Specifically, the solution is decomposed into two orthogonal components aligned with the sum $\mathrm{Im}\,({a}+{b})$ and the quaternion square root $\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}$. This explicit spanning set provides a direct geometric link between the equation parameters and the solution’s orientation. Such structural transparency is vital in engineering systems, where the physical interpretation of quaternionic variables relative to rotation axes is critical.
5. Solutions to the inhomogeneous singular Sylvester equation
In this section we present the conditions for solvability and the general solution to the inhomogeneous singular Sylvester equation. For similar quaternions a,b{a},{b} and a nonzero quaternion c{c}, the inhomogeneous singular Sylvester equation is given as
(5.1) ${a}{x}-{x}{b}={c}.$
If a,b{a},{b} are real numbers, (5.1) has no solution, since (ab)x=0({a}-{b}){x}=0 for all quaternions x{x}. Therefore, in the following theorems, we assume that a,b{a},{b} are nonreal quaternions. The solvability condition influences the structure of the solution.
Lemma 5.1.
Let a,b{a},{b} be similar quaternions and c{c} a nonzero quaternion. Then, (5.1) has a solution if and only if ${a}{c}={c}{b}^{*}$.
Proof.
By Corollary 4.1.1, we may assume that a,b{a},{b} are pure quaternions. Since $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$, we have
${a}{c}={a}({a}{x}-{x}{b})=-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}{x}-{a}{x}{b}=-{a}{x}{b}-{x}\mathopen{\lvert}{b}\mathclose{\rvert}^{2}=-({a}{x}-{x}{b}){b}=-{c}{b},$
and hence the assertion of the theorem. ∎
Lemma 5.2.
Let a,b{a},{b} be similar quaternions and c{c} a nonzero quaternion with ${a}{c}={c}{b}^{*}$. Then, for an arbitrary quaternion q{q}, the general solution to (5.1) is given as
(5.2) ${x}=({\mathrm{Im}\,{a}})\frac{{q}-{c}}{4\mathopen{\lvert}{\mathrm{Im}\,{a}}\mathclose{\rvert}^{2}}+\frac{{q}+{c}}{4\mathopen{\lvert}{\mathrm{Im}\,{b}}\mathclose{\rvert}^{2}}({\mathrm{Im}\,{b}}).$
Proof.
By Corollary 4.1.1, we may assume that a,b{a},{b} are pure quaternions. Substituting ${c}={a}{a}^{-1}{c}$ and ${c}={c}{b}^{-1}{b}$ into (5.1) yields
(5.3) $0={a}{x}-{x}{b}-\frac{1}{2}({a}{a}^{-1}{c}+{c}{b}^{-1}{b})={a}({x}+\frac{{a}{c}}{2\mathopen{\lvert}{a}\mathclose{\rvert}^{2}})-({x}-\frac{{c}{b}}{2\mathopen{\lvert}{b}\mathclose{\rvert}^{2}}){b}.$
Applying Lemma 4.2 to the homogeneous singular Sylvester equation (5.3), we have
(5.4) ${x}+\frac{{a}{c}}{2\mathopen{\lvert}{a}\mathclose{\rvert}^{2}}={a}{q}+{q}{b}\qquad\Longleftrightarrow\qquad{x}={a}{q}+{q}{b}-\frac{{a}{c}}{2\mathopen{\lvert}{a}\mathclose{\rvert}^{2}},$
for an arbitrary quaternion q{q}. Substituting ${a}{c}=\frac{1}{2}({a}{c}-{c}{b})$ into (5.4) yields (5.2). ∎
Theorem 5.3.
Let a,b{a},{b} be similar quaternions and c{c} a nonzero quaternion with ${a}{c}={c}{b}^{*}$. Then, for arbitrary real numbers $\lambda,\mu$ the general solution to (5.1) is given as
(5.5) ${x}=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\mathrm{Im}\,({a}+{b})+\frac{{c}}{4\mathopen{\lvert}{\mathrm{Im}\,{a}}\mathclose{\rvert}^{2}}\mathrm{Im}\,({b}-{a}).$
Proof.
From Lemma 4.2 and Theorem 4.3, for an arbitrary quaternion q{q} there exist real numbers $\lambda,\mu$ such that
$({\mathrm{Im}\,{a}}){q}+{q}({\mathrm{Im}\,{b}})=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\mathrm{Im}\,({a}+{b}),$
which applied to (5.2) yields (5.5). ∎
6. Conclusion
We have presented closed-form solutions to the homogeneous and inhomogeneous singular Sylvester equations in quaternions, together with a restatement of the corresponding solvability conditions. In addition, we derived an explicit formula for quaternion roots and showed how the closed-form solution naturally incorporates the quaternion square root. This structural connection provides a pathway for extending the methodology to a broader class of problems in quaternion analysis.
Data Availability Statement
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
Acknowledgement
The authors would like to thank the Federal Ministry of Research, Technology, and Space (BMFTR) for its support as part of the research program Communication Systems “Souverän. Digital. Vernetzt.”. Joint project 6G-life, project identification number: 16KIS2413K and the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/2 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of TUD Dresden University of Technology.